Find how many numbers of N atoms in 97.4 g NH4NO2


1
Expert's answer
2021-07-05T03:22:56-0400

molar mass of NH4NO2 = 64.0 g/mol

moles of NH4NO2 = mass / molar mass = (92.6/1000) g / 64.0 g/mol = 0.00145 mol

1 mole of NH4NO2 contains two moles of N atoms,

thus, moles of N atoms = moles of NH4NO2 x 2 = 0.00145 x 2 = 0.00290 mol

1 mole of N atoms contains 6.02 x 10^23 N atoms,

number of N atoms = 0.00290 x 6.02 x 10^23 = 1.75 x 10^21